# When Sharpening Becomes Collapse: Sampling Bias and Semantic Coupling in RL with Verifiable Rewards

### v1

@@ 2.3 Semantic Coupling
In particular, if $\lambda_{\min}\approx\lambda_{\max}\approx\lambda$ and $\rho_{\min}\approx\rho_{\max}\approx\rho$, $\mathcal{R}$ simplifies to $\left[\frac{N_{o^{\prime}}y_{k}}{\mathcal{A}_{\text{sum}}}\left(\frac{\lambda}{\rho}+G-1\right)-1\right].$ See Appendix F for proof.
Qualitatively, $\mathcal{R}$ measures how much more strongly a batch‑seen mode at $q^{\prime}$ is pushed compared to a semantically related but unseen mode. A large value of $\mathcal{R}$ means that the guaranteed logit increase of a batch‑seen mode at $q^{\prime}$ dominates the maximum possible increase of a batch‑unseen mode. In this regime, probability mass is pulled disproportionately toward the former, leading to over‑sharpening at $q^{\prime}$. From the explicit form, over‑sharpening is promoted when:
  • The update budget is concentrated on a single overlapping mode, i.e., $N_{o^{\prime}}y_{k}/\mathcal{A}_{\text{sum}}$ is large.
  • Features are highly specific, i.e., $\lambda/\rho$ is large, so updates align strongly with batch directions and leak little to other modes.
  • Effective batch reinforcement is strong, i.e., the batch size $G$ and the number of repeated samples of the same mode are large.
@@ 2.4 Empirical Illustration of Over‑Sharpening
Both Persian and Siamese are close to the Cat superclass while differing along their subclass‑specific dimensions. In this case, the input embedding plays the role of the query $q$, and the sampled label plays the role of the output $o$. We train this classifier with common RLVR algorithms, using a binary verifiable reward and SGD for optimization. We present only Persian inputs for 500 steps. At each step, we sample a group of $G$ labels from the current softmax policy and assign reward 1 if the sampled label is Cat or Persian, and 0 otherwise, i.e., both the superclass and subclass are treated as correct for Persian. We then compute advantages under various methods (raw reward, mean-shifted, and normalized) and update the policy accordingly. We track the probabilities $\pi(o=\text{Cat}|q=\text{Persian})$ and $\pi(o=\text{Persian}|q=\text{Persian})$ for different group sizes $G$, as shown in Figure 1.
Across all advantage estimators and group sizes, the model eventually collapses onto a single label (either Cat or Persian). Which label wins is random, determined by early sampling noise. Once one mode is sampled more often, it dominates subsequent updates. Increasing $G$ reduces the variance of early updates and thereby delays collapse, but does not prevent it. The choice of the advantage estimator primarily modulates the rate of collapse. Raw rewards lead to the most aggressive over‑sharpening, mean‑shifted advantages are slightly more conservative, and, normalized advantages slow the process further. This aligns with our theory that normalization (via $p^{+}(1-p^{+})/\sigma$) stabilizes $Z^{\prime}(q)$. We further evaluate advanced advantage estimators, including RLOO and Reinforce++ with baseline, and observe that while they modulate the convergence rate, the collapse persists. Moreover, momentum and adaptive optimizers (MI, AdamW) typically exacerbate over‑sharpening by accumulating past gradients and amplifying early sampling biases, pushing the policy more quickly into the collapsed regime. See Appendix G for these results.
We next study semantic coupling by measuring how training on Persian affects the held‑out Siamese subclass. We keep the training protocol unchanged but now monitor $\pi(o=\text{Siamese}\mid q=\text{Siamese})$ over training. Our theory predicts that the propagation of collapse is governed by $\mathcal{R}$ which grows with group size $G$ and with semantic coupling through the ratio $\lambda/\rho$. To modulate semantic coupling, we modify the Siamese embedding:
@@
progressively shifting mass from the Persian‑specific coordinate (first dimension) to the Siamese‑specific coordinate (third dimension). We refer to these as High, Mid, and Low similarity. As can be seen in Figure 2, we observe that the probability of correctly classifying Siamese is strongly squeezed over training. The effect is more severe when similarity is higher (larger shared features) and $G$ is larger. As we increase the overlap between Persian and Siamese features, the empirical ratio $\lambda/\rho$ estimated from the gradient Gram matrix also increases (details in Appendix G, which in turn raises $\mathcal{R}$. The observed faster collapse of Siamese accuracy thus matches the predicted growth in semantic over-sharpening. We highlight that this interaction between group size and semantic coupling is somewhat counterintuitive. Indeed, a larger $G$ is often expected to stabilize training, yet in our RLVR setting, it can worsen over‑sharpening and propagate collapse across semantically related classes.
@@ 3 Approach
We now introduce two complementary mechanisms designed to keep RLVR in moderate sharpening regime, including inverse-success advantage calibration (IAC) which adjusts how strongly each sampled trajectory contributes to the policy update, and distribution‑level calibration (DLC) which tunes the rollout distribution itself to prevent the policy from repeatedly sampling the same high‑probability patterns. Appendix H presents the full algorithmic pipeline (Algorithm 1), alongside visualizations that show the effectiveness of our methods in mitigating over-sharpening.
@@ 3.1 Inverse‑success Advantage Calibration (IAC)
Our first strategy targets the mechanism identified in Theorem 2.3: over‑sharpening at a query $q$ is driven by an inflated batch‑dependent partition function $Z^{\prime}(q)$, which uniformly suppresses the probabilities of all unsampled modes. Specifically, Theorem 2.3 gives $Z^{\prime}(q)\;\geq\;1+\frac{p^{+}(1-p^{+})}{\beta\sigma}\,\bigl[\pi_{\min}^{+}-\pi_{\max}^{-}\bigr].$ We find that standard advantage normalization cannot cancel the second term. When the reference model is already confident on positive samples $\pi_{\min}^{+}-\pi_{\max}^{-}>0$, the right‑hand side exceeds 1, causing $Z^{\prime}(q)>1$ and aggressive suppression of all unsampled (including correct) modes.
A natural solution would be to reweight advantages inversely proportional to $\pi_{\text{ref}}(\cdot)$ so that high‑probability easy positives contribute less. In detail, let $\tilde{A}_{k}=A(q,o_{k})/\pi_{\text{ref}}(o_{k})$. Then the lower bound becomes:
@@
In the last line, we use the normalization property $\sum_{k}N_{k}A_{k}=0$. In practice, however, we observe two issues when directly implementing this calibration:
@@
As a remedy, we only calibrate the advantages of positive samples, and we do so using an inverse‑success heuristic rather than the raw probabilities. For each positive sample, we define the following calibrated advantage:
$\tilde{A}(q,o_{k}^{+})\;=\;A(q,o_{k}^{+})\cdot(G-|S^{+}|)^{\alpha},\ \alpha\geq 0.$
Notice that we leave negative samples unmodified, i.e., $\tilde{A}(q,o_{k}^{-})=A(q,o_{k}^{-})$. This design has two key properties:
  • $(G-|S^{+}|)$ serves as a proxy for the difficulty of the query. When positives are rare, the factor is large. Thus, harder queries receive a larger effective weight.
  • The exponent $\alpha$ controls the calibration strength. $\alpha=0$ recovers the uncalibrated baseline, while larger $\alpha$ increases the relative weight of positives from hard queries and de‑emphasizes those from easy queries.
Intuitively, this inverse‑success calibration redistributes the advantage budget away from queries where the model is already confident and towards queries where success is rare, thereby preventing $Z^{\prime}(q)$ from becoming excessively large and mitigating over‑sharpening of already mastered modes. At the same time, by not amplifying negative samples, we avoid collapsing on partially correct or truncated trajectories.
@@ 3.2 Distribution-Level Calibration via Decoupled Sampling (DLC)
While inverse‑success calibration moderates how strongly each batch can push $Z^{\prime}(q)$, it does not address that once certain trajectories have been suppressed and their probabilities become very small, they almost never appear again in subsequent rollouts. This feedback loop, i.e., low probability leading to low sampling frequency, exacerbates sampling bias over time. Thus, our second mitigation is to calibrate the sampling distribution itself.
A vanilla idea would be to increase the sampling temperature of the policy, flattening the distribution. However, a global temperature treats all tokens equally, ignoring that different tokens learn at different speeds and have different difficulty. To address this, we propose a memory‑network–based distribution calibration that adapts exploration based on the policy’s rollout history.
In detail, we introduce an auxiliary memory network $m_{\phi}$ with logits $f_{\phi}(q,o)$, trained to approximate the empirical frequency of $(q,o)$ pairs encountered during training process. Concretely, after each training step, we update $m_{\phi}$ on the sampled trajectories (with a standard cross‑entropy loss) so that it outputs higher logits for frequently generated outputs and lower logits for rarely seen ones.
At sampling time, rather than drawing from the raw policy logits $f_{\theta}(q,\cdot)$, we construct a calibrated sampling distribution by subtracting the memory logits:
@@
Computational overhead and practical choice. Compared to inverse‑success calibration, the memory network is significantly more expensive. For instance, if $m_{\phi}$ shares the same architecture as the policy, then in principle we roughly double both training and rollout compute. In our implementation, the dominant additional cost comes from rollout. Although integrated with high-performance backends like vLLM, the absence of kernel-level support for synchronized multi-model sampling results in a $5\times$ rollout overhead. Given this overhead, our subsequent evaluations mainly use inverse‑success advantage calibration as the default mitigation, since it is lightweight. We treat the memory network as an optional, higher‑cost augmentation that can further improve performance when compute allows. Empirically, we observe that adding the memory network on top of inverse‑success calibration provides additional gains, but for most RLVR scenarios, the advantage‑level method alone strikes a better balance between effectiveness and efficiency.
Table 2: AVG@8 (%) performance of various methods across six benchmark datasets.
Model Method GSM8K Math500 LightEval AIME24 Minerva O-Bench Avg.
Qwen3-4B GRPO 93.72 85.58 70.61 34.58 43.06 48.85 62.73
RLOO 93.69 86.44 70.24 32.50 43.98 47.91 62.46
Reinforce++ 94.04 85.43 70.21 36.25 45.63 48.57 63.36
DrGRPO 94.10 83.56 69.54 35.42 45.43 47.83 62.65
IAC (Ours) 94.63 86.43 72.15 37.08 47.01 48.87 64.36
Qwen3-8B GRPO 95.59 85.81 71.12 35.42 47.20 49.79 64.16
RLOO 95.56 86.89 72.30 36.25 46.78 49.74 64.59
Reinforce++ 95.56 87.07 72.14 40.68 47.84 50.30 65.60
DrGRPO 95.35 86.11 71.56 34.58 47.38 49.87 64.14
IAC (Ours) 95.72 87.67 73.85 35.42 50.60 50.49 65.63
Qwen3-14B GRPO 95.96 89.42 73.83 53.45 50.32 55.28 69.71
RLOO 96.04 90.29 73.64 54.17 50.46 55.10 69.95
Reinforce++ 95.82 90.88 73.96 54.17 50.87 56.00 70.28
DrGRPO 95.96 90.37 73.85 54.85 50.69 55.35 70.18
IAC (Ours) 96.36 90.70 74.52 56.67 52.34 56.32 71.15
Table 3: PASS@8 (%) performance of various methods across six math benchmark datasets.
Model Method GSM8K Math500 LightEval AIME24 Minerva O-Bench Avg.
Qwen3-4B GRPO 96.74 91.94 75.43 47.37 54.54 56.44 70.41
RLOO 96.98 92.22 75.39 43.65 55.01 56.31 69.93
Reinforce++ 96.45 91.58 74.92 50.54 55.32 56.38 70.87
DrGRPO 96.43 89.90 73.87 47.02 54.60 55.08 69.48
IAC (Ours) 97.38 92.33 77.30 52.28 57.74 56.59 72.27
Qwen3-8B GRPO 96.51 91.39 75.19 52.36 55.65 56.88 71.33
RLOO 97.25 92.64 76.17 50.92 55.55 57.68 71.70
Reinforce++ 96.96 92.31 75.96 55.24 56.12 57.78 72.40
DrGRPO 97.11 91.57 75.36 52.25 55.41 57.26 71.49
IAC (Ours) 97.56 94.54 78.69 52.67 60.60 58.48 73.76
Qwen3-14B GRPO 97.53 93.75 76.54 70.44 59.40 60.87 76.42
RLOO 97.43 93.62 76.36 73.18 59.78 61.14 76.92
Reinforce++ 97.79 94.12 76.74 71.71 60.27 61.42 77.01
DrGRPO 97.53 94.27 76.57 71.83 61.03 60.96 77.03
IAC (Ours) 98.11 94.37 77.28 78.07 62.64 62.37 78.81
@@ 4.1 Experimental Setup
Models, datasets, and metrics. We evaluate our methods in the context of mathematical reasoning with verifiable rewards. We use the DeepScaleR (Luo et al., 2025) corpus as our primary RLVR training set, consisting of approximately 40.3k problem–answer pairs with checkable labels. To assess generalization, we report performance on six diverse benchmarks: GSM8K, AIME 2024, MATH 500, LightEval, Minerva (Lewkowycz et al., 2022), and O‑Bench (He et al., 2024). We conduct experiments with three backbone models, Qwen3‑4B, Qwen3-8B and Qwen3‑14B (Yang et al., 2025). As evaluation metrics, we report PASS@8 and AVG@8. PASS@8 measures the proportion of queries for which at least one of the eight responses is correct, and AVG@8 reflects the average proportion of correct responses per query.
Baselines. We compare our approaches against a range of RL baselines commonly used in LLM training, including GRPO (DeepSeek-AI, 2025), DrGRPO (Liu et al., 2025), RLOO (Kool et al., 2019), Reinforce++ with baseline (Hu et al., ), and DAPO (Yu et al., 2025). All baselines are run in the same setting with identical reward functions and sampling budgets for fair comparison.
Implementation details. All models are optimized with AdamW using a learning rate of $10^{-6}$. We employ a global batch size of 64 and sample $G=$ 8 responses per query. Training is run for 2000 steps, and we evaluate checkpoints every 50 steps, reporting the peak performance for each method. We employ VERL’s built-in recipes to standardize the implementation of baseline methods. For IAC, we use $\alpha=1$ by default. Due to the page limit, we leave sensitivity analysis ($\alpha$), the comparison with sample-filtering baselines, and the entropy evolution analysis in Appendix I. All experiments are conducted using 32 H200 GPUs.

### v2

@@ 3.3 Semantic Coupling
In particular, if $\lambda_{\min}\approx\lambda_{\max}\approx\lambda$ and $\rho_{\min}\approx\rho_{\max}\approx\rho$ (i.e., $\mathbf{K}_{qq}$ is nearly a diagonally dominant matrix with uniform diagonal and off‑diagonal entries), $\mathcal{R}$ simplifies to $\left[\frac{N_{o^{\prime}}y_{k}}{\mathcal{A}_{\text{sum}}}\left(\frac{\lambda}{\rho}+G-1\right)-1\right].$ See Appendix H for proof.
The ratio $\mathcal{R}$ captures the relative strength with which a batch‑seen mode at $q^{\prime}$ is reinforced compared to a semantically related but batch‑unseen mode. A large $\mathcal{R}$ indicates that the guaranteed increase for the seen mode significantly outweighs the maximum possible increase for the unseen mode, leading to a disproportionate redistribution of probability mass and thus over‑sharpening at $q^{\prime}$. From the explicit form, over‑sharpening is promoted when:
  • A large value of $N_{o^{\prime}}y_{i}/\mathcal{A}_{\text{sum}}$ means that most of the update budget is focused on the particular mode that overlaps between $q$ and $q^{\prime}$.
  • Large $\lambda/\rho$ corresponds to highly specific features: gradient updates are tightly aligned with individual modes and leak only weakly to other modes. In this regime, improvement on $q$ reinforces a particular mode at $q^{\prime}$ rather than broadly supporting all semantically consistent solutions.
  • A larger effective batch size $G$ amplifies the effect.
@@ 3.4 Empirical Illustration of Over‑Sharpening
Persian and Siamese align with the Cat superclass but differ in subclass dimensions. Treating inputs as $q$ and labels as $o$, we train this classifier on Persian inputs for 500 steps using SGD. At each step, we sample $G$ outputs labels from the current softmax policy and assign reward 1 if the sampled label is Cat or Persian, and 0 otherwise, i.e., both the superclass and subclass are treated as correct for Persian. We update the policy using various advantage estimators (raw, mean-shifted, normalized) and track the probabilities $\pi(o=\text{Cat}|q=\text{Persian})$ and $\pi(o=\text{Persian}|q=\text{Persian})$ for different group sizes $G$, as shown in Figure 1.
Across all settings, the model eventually collapses onto a single label (either Cat or Persian, driven randomly by early sampling noise). Once one mode is sampled more often, it dominates subsequent updates. Increasing $G$ can delay collapse, but does not prevent this collapse. The choice of the advantage estimator primarily modulates the rate of collapse. Raw rewards are most aggressive, while normalized advantages are the slowest, aligning with our theory on $Z^{\prime}(q)$ stabilization. We also evaluate advanced advantage estimators, including RLOO and Reinforce++ with baseline, and observe that the collapse persists. Moreover, momentum and adaptive optimizers (MI, AdamW) typically exacerbate over‑sharpening by accumulating past gradients and amplifying early sampling biases, pushing the policy more quickly into the collapsed regime. See Appendix I for these results.
We next study semantic coupling by monitoring the held-out Siamese subclass ($\pi(o=\text{Siamese}\mid q=\text{Siamese})$) while training on Persian. Our theory shows that the propagation of collapse is governed by the interplay between the transfer intensity $\eta$ (Theorem 3.4), and the relative suppression ratio $\mathcal{R}$. Notably, $\mathcal{R}$ scales with $G$. To test this, we vary Siamese similarity (High, Mid, Low) by progressively shifting embedding mass from Persian-specific (dim 1) to Siamese-specific (dim 3) coordinates:
@@
As shown in Figure 2, Siamese accuracy is strongly squeezed during training. This effect worsens with higher similarity, confirming that increased $\eta$ (high similarity) accelerates collapse. Similarly, larger $G$ exacerbates over-sharpening.
@@ 4 Approach
We introduce two complementary mechanisms to keep RLVR in a moderate sharpening regime. Inverse-success advantage calibration (IAC) adjusts how strongly each sampled trajectory contributes to the policy update. Distribution-level calibration (DLC) tunes the rollout distribution to prevent the policy from repeatedly sampling the same high-probability patterns. Appendix J details the full pipeline (Algorithm 1) and visualizations demonstrating their effectiveness in mitigating over-sharpening via the Softmax case used in Section 3.4.
@@ 4.1 Inverse‑success Advantage Calibration (IAC)
Idealized inverse‑probability reweighting. A natural solution would be to reweight advantages inversely proportional to $\pi_{\text{ref}}(\cdot)$, so that the reference‑probability factor in the bound cancels out. In detail, let $\tilde{A}_{k}=A(q,o_{k})/\pi_{\text{ref}}(o_{k})$. Then the lower bound becomes:
@@
In the last line, we use the normalization property $\sum_{k}N_{k}A_{k}=0$. However, this ideal solution is not directly usable in large‑scale RLVR for two reasons:
@@
Inverse-success advantage calibration. Motivated by the above, we retain the spirit of inverse reweighting but replace raw probabilities by a coarse, query‑level difficulty proxy. We only calibrate positive advantages and leave negative ones untouched. For each query $q$, let $S^{+}$ be the set of positive trajectories in the batch and $G$ the total number of trajectories. We define the calibrated advantages:
$\tilde{A}(q,o_{k}^{+})$ $=A(q,o_{k}^{+})\cdot(G-|S^{+}|)^{\alpha},\quad\alpha\geq 0,$ (4)
$\tilde{A}(q,o_{k}^{-})$ $=A(q,o_{k}^{-}),$ (5)
where $o_{k}^{+}$ and $o_{k}^{-}$ denote positive and negative trajectories, respectively. This simple heuristic enjoys two desirable properties: This design has two key properties. (1) $(G-|S^{+}|)$ serves as a difficulty proxy, assigning larger weights to harder queries where positive samples are rare. (2) The exponent $\alpha$ governs the overall calibration strength. $\alpha=0$ recovers the baseline, while larger values de-emphasize easy queries. Intuitively, IAC redistributes the advantage budget from easy queries to difficult ones, preventing excessive growth of $Z^{\prime}(q)$ and mitigating over-sharpening. By keeping negative advantages unchanged, we can avoid destabilizing training on partially correct or truncated trajectories.
@@ 4.2 Distribution-Level Calibration via Decoupled Sampling (DLC)
While IAC moderates how strongly each batch can inflate $Z^{\prime}(q)$, it does not address that once certain trajectories have been suppressed and their probabilities become very small, they almost never appear again in subsequent rollouts. To mitigate this, we suggest calibrating the sampling distribution itself.
A vanilla approach would be to increase the global sampling temperature, thereby flattening the policy distribution. However, a global temperature treats all tokens equally, some modes are already well‑explored and saturated, while others remain under‑explored. We instead propose a distribution-Level dalibration (DLC) based on a memory network that tracks the empirical rollout distribution.
DLC introduces an auxiliary memory network $m_{\phi}$ with logit function $f_{\phi}(q,o)$, intended to approximate the empirical frequency of $(q,o)$ pairs generated during training. At each training step, we update $m_{\phi}$ on the sampled trajectories (with a standard cross‑entropy loss) so that it outputs higher logits for frequently generated outputs and lower logits for rarely seen ones.
During rollout, DLC constructs a calibrated sampling distribution by subtracting the memory logits from the policy logits as follows:
@@
Computational overhead and practical choice. Compared to IAC, DLC introduces non‑trivial computational overhead because it requires both policy and memory forward passes during rollout, and training updates for $m_{\phi}$. If $m_{\phi}$ shares the same architecture as the policy, the theoretical cost roughly doubles. In our implementation, the dominant additional cost comes from rollout. Although integrated with high-performance rollout backends like vLLM, the lack of kernel‑level support for synchronized multi‑model sampling leads to approximately $5\times$ higher rollout latency. Given this overhead, our default configuration in large‑scale experiments employs IAC alone, which is compute‑efficient and already substantially mitigates over‑sharpening. We treat DLC as an optional augmentation that can be enabled when compute permits. Empirically, adding DLC on top of IAC yields further improvements, especially in regimes where over‑exploitation of a few modes is particularly pronounced.
Table 2: Comparison of AVG@8 and PASS@8 performance across six benchmark datasets.
Model Method AVG@8 (%) PASS@8 (%)
GSM8K Math500 LightEval AIME24 Minerva O-Bench Avg. GSM8K Math500 LightEval AIME24 Minerva O-Bench Avg.
Qwen3-4B GRPO 93.72 85.58 70.61 34.58 43.06 48.85 62.73 96.74 91.94 75.43 47.37 54.54 56.44 70.41
RLOO 93.69 86.44 70.24 32.50 43.98 47.91 62.46 96.98 92.22 75.39 43.65 55.01 56.31 69.93
Reinforce++ 94.04 85.43 70.21 36.25 45.63 48.57 63.36 96.45 91.58 74.92 50.54 55.32 56.38 70.87
DrGRPO 94.10 83.56 69.54 35.42 45.43 47.83 62.65 96.43 89.90 73.87 47.02 54.60 55.08 69.48
IAC (Ours) 94.63 86.43 72.15 37.08 47.01 48.87 64.36 97.38 92.33 77.30 52.28 57.74 56.59 72.27
Qwen3-8B GRPO 95.59 85.81 71.12 35.42 47.20 49.79 64.16 96.51 91.39 75.19 52.36 55.65 56.88 71.33
RLOO 95.56 86.89 72.30 36.25 46.78 49.74 64.59 97.25 92.64 76.17 50.92 55.55 57.68 71.70
Reinforce++ 95.56 87.07 72.14 40.68 47.84 50.30 65.60 96.96 92.31 75.96 55.24 56.12 57.78 72.40
DrGRPO 95.35 86.11 71.56 34.58 47.38 49.87 64.14 97.11 91.57 75.36 52.25 55.41 57.26 71.49
IAC (Ours) 95.72 87.67 73.85 35.42 50.60 50.49 65.63 97.56 94.54 78.69 52.67 60.60 58.48 73.76
Qwen3-14B GRPO 95.96 89.42 73.83 53.45 50.32 55.28 69.71 97.53 93.75 76.54 70.44 59.40 60.87 76.42
RLOO 96.04 90.29 73.64 54.17 50.46 55.10 69.95 97.43 93.62 76.36 73.18 59.78 61.14 76.92
Reinforce++ 95.82 90.88 73.96 54.17 50.87 56.00 70.28 97.79 94.12 76.74 71.71 60.27 61.42 77.01
DrGRPO 95.96 90.37 73.85 54.85 50.69 55.35 70.18 97.53 94.27 76.57 71.83 61.03 60.96 77.03
IAC (Ours) 96.36 90.70 74.52 56.67 52.34 56.32 71.15 98.11 94.37 77.28 78.07 62.64 62.37 78.81
@@ 5.1 Experimental Setup
Models, datasets, and metrics. We evaluate our methods in mathematical reasoning with verifiable rewards. We use the DeepScaleR (Luo et al., 2025) corpus as our primary RLVR training set, consisting of approximately 40.3k problem–answer pairs. To assess generalization, we report performance on six diverse benchmarks: GSM8K, AIME 2024, MATH 500, LightEval, Minerva (Lewkowycz et al., 2022), and O‑Bench (He et al., 2024). We conduct experiments with three backbone models, Qwen3‑4B, Qwen3-8B and Qwen3‑14B (Yang et al., 2025). As evaluation metrics, we report PASS@8 and AVG@8. PASS@8 measures the proportion of queries for which at least one of the eight responses is correct, and AVG@8 reflects the average proportion of correct responses per query. Appendix K provides evaluation results on different models and datasets.
Baselines. We compare our approaches against a range of RL baselines commonly used in LLM training, including GRPO (DeepSeek-AI, 2025), DrGRPO (Liu et al., 2025), RLOO (Kool et al., 2019), Reinforce++ with baseline (Hu et al., ), and DAPO (Yu et al., 2025). We also compare IAC with sample-filtering-based methods (Bae et al., 2025) in Appendix K. All baselines are run in the same setting with identical reward functions and sampling budgets for fair comparison.
Implementation details. All models (including the memory network) are optimized with AdamW using a learning rate of $10^{-6}$. We employ a global batch size of 64 and sample $G=$ 8 responses per query. Training is run for 2000 steps, and we evaluate checkpoints every 50 steps, reporting the peak performance for each method. We employ VERL’s built-in recipes to standardize the implementation of baseline methods. For IAC, we use $\alpha=1$ by default. Due to the page limit, we leave sensitivity analysis ($\alpha$), the comparison with sample-filtering baselines, and the entropy evolution analysis in Appendix K. All experiments are conducted using 32 H200 GPUs.